OpenStudy (anonymous):

CHECK ANSWER PLEASE ! MEDAL WILL BE GIVEN ! Which ordered pair makes the equation true? 2x – 5y = –10 (5, 4) (–10, 4) (2, 5) (20, 4) C ? ?

OpenStudy (anonymous):

please help (:

OpenStudy (therealmeeeee):

5, 4) (–10, 4) (2, 5) (20, 4) The first number = x The second number = y Plug in the numbers for x and y and simplify I'll do the 3rd one for an example 2(2)-5(5)=-10 4 -25=-10 -21 = -10 Well, -21 does not equal -10, so that's not the right answer

OpenStudy (therealmeeeee):

its B

OpenStudy (anonymous):

aww thanks i really needed that ... @TheRealMeeeee

OpenStudy (therealmeeeee):

np